What is the equation of the line passing through the points (2, –1) and (5, –10) in slope-intercept form?
Found 2 solutions by Cromlix, fractalier:
Answer by Cromlix(4381)

Hi there,
Slope = y2 - y1/x2 - x1 = -10 -(-1)/5 - 2
slope = -9/3 = -3
Using line equation
y - b = m(x - a)
where m = -3 and (a,b) = (2, -1)
y - (-1) = -3(x - 2)
y + 1 = -3x + 6
y = -3x + 6 - 1
y = -3x + 5
Hope this helps :-)

Answer by fractalier(6550)

Couple of ways to do this...let's find the slope first...
slope m = (-10 - (-1)) / (5 - 2) = -9/3 = -3
Now plug into y = mx + b to find b...like this
-1 = -3(2) + b
b = 5
so that our line is
y = -3x + 5